# Statistical Inference for Policy Evaluation with Temporal Difference Learning

### v3

@@ D.1 Online algorithm to construct the variance estimator
2:Initialize $\bm{\theta}=\bm{0}$, $\overline{\bm{\theta}}=\bm{0}$, $\overline{\bm{A}}=\bm{0}$, $\overline{\bm{AA}}=\bm{0}$, $\overline{\bm{Ab}}=\bm{0}$, $\overline{\bm{bb}}=\bm{0}$
3:for $t=0:(T-1)$ do
4:     Sample $s\sim\mu$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s)$
5:     Compute:
$\bm{A}\leftarrow\bm{\phi}(s)[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})],\quad\bm{b}\leftarrow r(s)\bm{\phi}(s)$
6:     Perform TD update:
$\bm{\theta}\leftarrow\bm{\theta}-\eta_{0}t^{-\alpha}(\bm{A}\bm{\theta}-\bm{b})$
7:     Perform online updates:
$\overline{\bm{\theta}}\leftarrow\overline{\bm{\theta}}+\frac{1}{t}(\bm{\theta}-\overline{\bm{\theta}})$
$\overline{\bm{A}}\leftarrow\overline{\bm{A}}+\frac{1}{t}(\bm{A}-\overline{\bm{A}})$
$\overline{\bm{AA}}\leftarrow\overline{\bm{AA}}+\frac{1}{t}(\bm{A}\otimes\bm{A}-\overline{\bm{AA}})$
$\overline{\bm{Ab}}\leftarrow\overline{\bm{Ab}}+\frac{1}{t}(\bm{A}\otimes\bm{b}+\bm{b}\otimes\bm{A}-\overline{\bm{Ab}})$
$\overline{\bm{bb}}\leftarrow\overline{\bm{bb}}+\frac{1}{t}(\bm{b}\otimes\bm{b}-\overline{\bm{bb}})$
8:end for
9:Construct $\widehat{\bm{\Gamma}}$:
$\widehat{\bm{\Gamma}}\leftarrow\mathsf{reshape}(\overline{\bm{AA}}(\overline{\bm{\theta}}\otimes\overline{\bm{\theta}})-2\overline{\bm{Ab}}(\overline{\bm{\theta}})+\overline{\bm{bb}},(d\times d))$
10:Construct the variance estimator:
$\widehat{\bm{\Lambda}}\leftarrow\overline{\bm{A}}^{-1}\widehat{\bm{\Gamma}}\overline{\bm{A}}^{-\top}$
11:Averaged TD estimator $\overline{\bm{\theta}}$, covariance matrix estimator $\widehat{\bm{\Lambda}}$
The basic properties of Kronecker products indicate
$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}\right)=(\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T});$
$\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}\right)=(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T};$
$\textbf{vec}\left(\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}\right)=(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T};$
$\textbf{vec}\left(\bm{b}_{t}\bm{b}_{t}^{\top}\right)=\bm{b}_{t}\otimes\bm{b}_{t}.$
Therefore, the vectorization of $\widehat{\bm{\Gamma}}_{T}$ can be represented as
$\textbf{vec}(\widehat{\bm{\Gamma}}_{T})$ $=\frac{1}{T}\sum_{t=1}^{T}\textbf{vec}\left(\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{\theta}}_{T}^{\top}\bm{A}_{t}-\bm{A}_{t}\overline{\bm{\theta}}_{T}\overline{\bm{b}}_{t}^{\top}-\bm{b}_{t}\bm{\overline{\theta}}_{T}^{\top}\bm{A}_{t}^{\top}+\bm{b}_{t}\bm{b}_{t}^{\top}\right)$
$=\frac{1}{T}\sum_{t=1}^{T}\left((\bm{A}_{t}\otimes\bm{A}_{t})(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-(\bm{A}_{t}\otimes\bm{b}_{t})\overline{\bm{\theta}}_{T}-(\bm{b}_{t}\otimes\bm{A}_{t})\overline{\bm{\theta}}_{T}+\bm{b}_{t}\otimes\bm{b}_{t}\right)$
=($=\underset{I_{1}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}\bm{A}_{t}\otimes\bm{A}_{t}\right)}}(\overline{\bm{\theta}}_{T}\otimes\overline{\bm{\theta}}_{T})-\underset{I_{2}}{\underbrace{\left(\frac{1}{T}\sum_{t=1}^{T}[\bm{A}_{t}\otimes\bm{b}_{t}+\bm{b}_{t}\otimes\bm{A}_{t}]\right)}}\overline{\bm{\theta}}_{T}$
$+\underset{I_{3}}{\underbrace{\frac{1}{T}\sum_{t=1}^{T}\bm{b}_{t}\otimes\bm{b}_{t}}},$
where $I_{1}$, $I_{2}$ and $I_{3}$ are sample means and can hence be calculated using online updates. Based on this, we propose Algorithm 1 for the calculation $\widehat{\bm{\Lambda}}_{T}$ in an online manner.
Since Algorithm 1 iterates from $t=0$ to $t=T-1$ in a single loop and does not store the samples for each iteration, it has a space complexity is $O(Td^{4})$ and a space complexity of $O(d^{4})$. Here, the $d^{4}$ term stems from calculating the Kronecker product $\bm{A}\otimes\bm{A}$.
@@ D.2 The Markov Decision Process and the feature map
=$=\left\{\begin{array}[]{lcc}q_{s}\mathds{1}(s^{\prime}=s)+\frac{1-q_{s}}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)&\text{for}&s<d;\\[5.69046pt] \frac{\gamma}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)+\frac{1-q_{s^{\prime}}}{d-1}\mathds{1}(s^{\prime}<d)&\text{for}&s\geq d.\end{array}\right.$ (116)
@@
From the expressions above, we remark that, the values of $q$ and $V^{\star}(s)$ with $s\geq d$ are fixed for all $\bm{q}\in\mathcal{Q}$ which is ensured by the construction (117). It is shown in Theorem 2 of Li et al. (2024) that the sequence $\{\mathcal{M}_{\bm{q}}\}_{q\in\mathcal{Q}}$ is hard to distinguish for a considerably large set $\mathcal{Q}$.
@@ D.3 Parameter specification
In our experiments, we fix the size of the state space to $\|\mathcal{S}\|=10$, the discount factor to be $\gamma=0.2$, and let $\varepsilon=0.01$. For the TD learning algorithm, we set the initial stepsize to be $\eta_{0}=5$, and total number of iterations to be $T=10^{5}$. Data is recorded every $100$ iterations. In order to show the role played by the stepsize decaying speed $\alpha$, we fix $d=3$ and $\alpha$ iterate through the group $(\frac{1}{2},\frac{2}{3},\frac{3}{4},\frac{4}{5})$; and in order to illustrate the influence of the dimension of the feature map, we fix $\alpha=\frac{2}{3}$ and let $d$ iterate through the group of $(3,5,7,9)$. For every given $d$, we set $\bm{q}$ to be
@@
    In order to illustrate the efficiency of our variance estimator $\widehat{\bm{\Lambda}}_{T}$, we measure the averaged Frobenius norm of the difference between $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ among $100$ independent trials. As is indicated by A.14, this is a direct proxy for the TV distance between the zero-mean Gaussian distributions with the two corresponding variance matrices.
@@
    In order to validate our statistical inference procedure, we run Algorithm 1 on the MDP with $d=3$ using $\alpha=\frac{2}{3}$, construct both individual and simultaneous confidence intervals for $\bm{\theta}^{\star}$, and measure the coverage rates of these confidence intervals in $1000$ independent trials. Details to follow in section D.4.4.
@@ D.4.1 High-probability convergence guarantee for the TD estimator
Figure 3 shows the decay of the $95\%$ quantile of the $L_{2}$ norm of $\overline{\bm{\Delta}}_{T}$; the slope of this log-log plot verifies our theoretical claim in Theorem 3.1 that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of $O(T^{-\frac{1}{2}})$ with high probability. Plot (a) shows that the convergence rate is similar for different choices of $\alpha>\frac{1}{2}$ when $T$ is large, while plot(b) shows that the convergence is slower when $d$ is larger. Both are in accordance with our conclusion in Theorem (1) that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of
@@ D.4.2 Berry-Esseen bound
Figure 4 illustrates how the finite-sample distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ converges to its asymptotic Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$. Notably, plot (a), which is a replication of Figure 2 in Section 4, indicates that the convergence is slower when $\alpha=\frac{1}{2}$. This directly contradicts the suggestion by Samsonov et al. (2024) to choose $\alpha=\frac{1}{2}$ in order to yield a faster in-distribution convergence of the TD estimation error. In fact, with $d=3$ fixed and other choices of $\alpha$, the convergence rate are similar when $T$ is large. Meanwhile, plot (b) shows the convergence generally is slower when $d$ is larger.
@@ D.4.3 Estimator of the covariance matrix
Figure 5 illustrates the rate by which our variance estimator $\widehat{\bm{\Lambda}}_{T}$ converges to the asymptotic variance $\bm{\Lambda}^{\star}$. Plot (a) shows that the convergence rate is similar for different choices of $\alpha$ when the number of interations $T$ is sufficiently large; plot(b) indicates that the convergence is slower when $d$ is larger, by the nearly parallel curves in the log-log plot demonstrate that the convergence rate with respect to $T$ is similar among different choices of $d$.
@@ D.4.4 Confidence intervals
Figure 7 shows the coverage rates of the independent and simultaneous confidence intervals, defined as the percentage of times when the optimal coefficient $\bm{\theta}^{\star}$ (or its entries) are covered by the confidence intervals in the $1000$ independent trials. As is indicated by the plot, the coverage rates all approach the target confidence level of $1-\delta=95\%$ when $T$ grows large.

### v4

@@ D.1 Online algorithm to construct the variance estimator
The code for our experiments can be found at https://www.dropbox.com/scl/fo/u28832upzxu7l24ewb0ar/AKWC7X68R7u3UEmgZ8PY5k0?rlkey=znzky05vh7q8gwc6bdmzhq5ms&st=4zsj1f02&dl=0. Experiments were performed on a personal laptop (MacOS). No external GPU clusters or cloud resources were used.
2:Initialize $\bm{\theta}=\bm{0}$, $\overline{\bm{\theta}}=\bm{0}$, $\overline{\bm{A}}=\bm{0}$, $\widehat{\bm{\Gamma}}=\bm{0}$.
3:for $t=1,\dots,T$ do
4: Sample $s\sim\mu$, $s^{\prime}\sim\mathcal{P}(\cdot\mid s)$.
5: Compute: $\bm{A}\leftarrow\bm{\phi}(s)[\bm{\phi}(s)-\gamma\bm{\phi}(s^{\prime})],\ \bm{b}\leftarrow r(s)\,\bm{\phi}(s)$.
6: Perform TD update: $\bm{\theta}\leftarrow\bm{\theta}-\eta_{0}t^{-\alpha}(\bm{A}\bm{\theta}-\bm{b})$.
7: Perform online updates:
[…]
11:end for
12:Construct the variance estimator:
[…]
14:Averaged TD estimator $\overline{\bm{\theta}}$, covariance matrix estimator $\widehat{\bm{\Lambda}}$.
@@
It is easy to verify that by definition (17), the estimators $\overline{\bm{A}}_{T}$ and $\widehat{\bm{\Gamma}}_{T}$ can be updated by
$\overline{\bm{A}}_{T}=\overline{\bm{A}}_{T-1}+\frac{1}{T}(\bm{A}_{T}-\overline{\bm{A}}_{T-1}),\quad\text{and}$
$\widehat{\bm{\Gamma}}_{T}=\widehat{\bm{\Gamma}}_{T-1}+\frac{2}{T+1}\left((\bm{A}_{T}\overline{\bm{\theta}}_{T}-\bm{b}_{T})(\bm{A}_{T}\overline{\bm{\theta}}_{T}-\bm{b}_{T})^{\top}-\widehat{\bm{\Gamma}}_{T-1}\right).$
Algorithm 1 reveals the complete process of generating both the averaged TD estimator $\overline{\bm{\theta}}_{T}$ and the variance estimator $\widehat{\bm{\Lambda}}_{T}$. Since Algorithm 1 iterates from $t=0$ to t=T1t=T-1 in a single loop and does not store the samples for each iteration, it has a space complexity is $O(Td^{2}+d^{3})$ and a space complexity of $O(d^{2})$. Here, the $d^{3}$ term stems from calculating matrix inverse.
@@ D.2 The Markov Decision Process and the feature map
$=\left\{\begin{array}[]{lcc}q_{s}\mathds{1}(s^{\prime}=s)+\frac{1-q_{s}}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)&\text{for}&s<d;\\[5.69046pt] \frac{\gamma}{|\mathcal{S}|-d+1}\mathds{1}(s^{\prime}\geq d)+\frac{1-q_{s^{\prime}}}{d-1}\mathds{1}(s^{\prime}<d)&\text{for}&s\geq d.\end{array}\right$. (118)
@@
From the expressions above, we remark that, the values of $q$ and $V^{\star}(s)$ with $s\geq d$ are fixed for all $\bm{q}\in\mathcal{Q}$ which is ensured by the construction (119). It is shown in Theorem 2 of Li et al., 2024b that the sequence $\{\mathcal{M}_{\bm{q}}\}_{q\in\mathcal{Q}}$ is hard to distinguish for a considerably large set $\mathcal{Q}$.
@@ D.3 Parameter specification
In our experiments, we fix the size of the state space to $\|\mathcal{S}\|=10$, the discount factor to be $\gamma=0.2$, and let $\varepsilon=0.01$. For the TD learning algorithm, we set the initial stepsize to be $\eta_{0}=5$, and total number of iterations to be $T=2^{20}$. Data is recorded at every power of 2, so that the log plot has constant intervals on the x-axis. In order to show the role played by the stepsize decaying speed $\alpha$, we fix $d=3$ and $\alpha$ iterate through the group $(\frac{1}{2},\frac{2}{3},\frac{3}{4},1$; and in order to illustrate the influence of the dimension of the feature map, we fix $\alpha=\frac{2}{3}$ and let $d$ iterate through the group of $(3,5,7,9)$. For every given $d$, we set $\bm{q}$ to be
@@
    In order to illustrate the efficiency of our variance estimator $\widehat{\bm{\Lambda}}_{T}$, we measure the averaged Frobenius norm of the difference between $\widehat{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$ among $10,000$ independent trials. As is indicated by A.14, this is a direct proxy for the TV distance between the zero-mean Gaussian distributions with the two corresponding variance matrices.
@@
    In order to validate our statistical inference procedure, we run Algorithm 1 on the MDP with $d=3$ using $\alpha=\frac{2}{3}$, construct both individual and simultaneous confidence intervals for $\bm{\theta}^{\star}$, and measure the coverage rates of these confidence intervals in $10,000$ independent trials. Details to follow in section D.4.4.
@@ D.4.1 High-probability convergence guarantee for the TD estimator
Figure 3 shows the decay of the $95\%$ quantile of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$; the slope of this log-log plot verifies our theoretical claim in Theorem 3.1 that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of $O(T^{-\frac{1}{2}})$ with high probability.
Plot (a) illustrates the behavior of the estimation error under different choices of $\alpha$. As a reference, the black dotted line represents the $95\%$ quantile of $\{\bm{Z}_{T}\}$ where $\bm{Z}_{T}\sim\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}/T)$. Our theory predicts that when $T$ grows large, this should be the asymptotic line for the empirical $95\%$ quantiles when $\alpha\in(\frac{1}{2},1)$. This is exactly what we observe that $\alpha=\frac{2}{3}$ and $\alpha=\frac{3}{4}$. On the contrary, when $\alpha=1$, the empirical quantiles does not converge to the asymptotic quantiles. This is also predicted by our theory that since $\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}=\bm{O}(T^{\alpha-1})$ (which does not converge when $\alpha=1$), the variance of $\overline{\bm{\Delta}}_{T}$ will not converge to $\bm{\Lambda}^{\star}$ when $\alpha=1$. This fact is also reflected in the $O(T^{\frac{\alpha}{2}-1})$ term in the bound of Theorem 3.1, which when $\alpha=1$, converges by $O(T^{-\frac{1}{2}})$, the same rate as the leading term. As for the case of $\alpha=\frac{1}{2}$, we observe that when $T$ is small, $\|\overline{\bm{\Delta}}_{T}\|_{2}$ is relatively large; it is not until a very large number of iterations (i.e., $T>10^{5}$), that the empirical and asymptotic lines begin to overlap. This phenomenon is also predicted by the $O(T^{-\frac{\alpha+1}{2}})$ term in the bound of Theorem 3.1, which converges slower when $\alpha=\frac{1}{2}$. A more detailed explanation follows in Appendix D.5.
Meanwhile, plot(b) shows that, fixing $\alpha=\frac{2}{3}$, the convergence is slower when $d$ is larger. Both are in accordance with our conclusion in Theorem (1) that $\|\overline{\bm{\Delta}}_{T}\|_{2}$ converges by the rate of
@@ D.4.2 Berry-Esseen bound
Figure 4 illustrates how the finite-sample distribution of $\sqrt{T}\overline{\bm{\Delta}}_{T}$ converges to its asymptotic Gaussian distribution $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$. For ease of computation, instead of the convex distance we evaluated the distance
$\max_{1\leq j\leq d}\sup_{x\in\mathbb{R}}\left|\widehat{\mathbb{P}}(\sqrt{T}(\overline{\bm{\Delta}}_{T})_{j}\leq x)-\mathbb{P}_{\bm{z}}((\bm{\Lambda}^{\star\frac{1}{2}}\bm{z})_{j}\leq x)\right|,$ (124)
where $\widehat{\mathbb{P}}$ refers to the empirical distribution in our $10,000$ trials, and $\bm{z}$ is the $d$-dimensional standard Gaussian.
Notably, plot (a) indicates that the convergence is slower when $\alpha=\frac{1}{2}$, compared to the choices of $\alpha=\frac{2}{3}$ and $\alpha=\frac{3}{4}$. This directly contradicts the suggestion by Samsonov et al., 2024a to choose $\alpha=\frac{1}{2}$ in order to yield a faster in-distribution convergence of the TD estimation error. On the other hand, when $\alpha=1$, the empirical distribution does not converge to $\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$; that verifies our claim in Appendix B.2.3 that the $O(T^{\alpha-1})$ term in the Berry-Esseen bound is unavoidable.
Meanwhile, plot (b) shows the convergence generally is slower when $d$ is larger.
@@ D.4.3 Estimator of the covariance matrix
Figure 5 illustrates the rate by which our variance estimator $\widehat{\bm{\Lambda}}_{T}$ converges to the asymptotic variance $\bm{\Lambda}^{\star}$. Plot (a) shows that the convergence rate is similar for different choices of $\alpha$ when the number of iterations $T$ is sufficiently large; plot(b) indicates that the convergence is slower when $d$ is larger, by the nearly parallel curves in the log-log plot demonstrate that the convergence rate with respect to $T$ is similar among different choices of $d$.
@@ D.4.4 Confidence intervals
Figure 7 shows the coverage rates of the individual and simultaneous confidence intervals, defined as the percentage of times when the optimal coefficient $\bm{\theta}^{\star}$ (or its entries) are covered by the confidence intervals in the $10,000$ independent trials. As is indicated by the plot, the coverage rates all approach the target confidence level of $1-\delta=95\%$ when $T$ grows large.
@@
D.5 Decomposition of TD estimation error
In this section, we aim to further explore the emergence of the TD estimation error $\overline{\bm{\Delta}}_{T}$ in numerical simulations, in order to verify our key claims in the proofs of our theoretical results. Specifically, we aim to explore the role played by the choice of $\alpha$ through decomposing the TD estimation error. Recall that throughout our proofs, we repetitively used the decomposition
$\overline{\bm{\Delta}}_{T}=\underset{I_{1}}{\underbrace{\frac{1}{T\eta_{0}}\bm{Q}_{0}\bm{\Delta}_{0}}}-\underset{I_{2}}{\underbrace{\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})}}-\underset{I_{3}}{\underbrace{\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}}},$
where the matrices $\{\bm{Q}_{i}\}_{0\leq i\leq T}$ are defined by (53), and therefore decided by the matrix $\bm{A}$ as well as the choice of stepsizes $\{\eta_{t}\}_{1\leq t\leq T}$. Of the three terms on the right-hand-side:
  1. 1.
    The term $I_{1}$ is related only to the initial estimator $\bm{\theta}_{0}$ through $\bm{\Delta}_{0}$ and not to the samples $\{(\bm{A}_{t},\bm{b}_{t})\}_{1\leq t\leq T}$. Therefore, we refer to this term as the initialization error. This term is deterministic, and according to our analysis, converges by the rate of
    $\left\|\frac{1}{T\eta_{0}}\bm{Q}_{0}\bm{\Delta}_{0}\right\|_{2}\lesssim O(T^{-1}).$
  2. 2.
    The term $I_{2}$ is related only to the samples $\{(\bm{A}_{t},\bm{b}_{t})\}_{1\leq t\leq T}$ and not to any of the iterations $\{\bm{\theta}_{t}\}_{0\leq t\leq T}$. Therefore, we refer to this term as the sampling error. According to our analysis, this term corresponds to the asymptotic error:
    $\frac{1}{\sqrt{T}}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\xrightarrow{d}\mathcal{N}(\bm{0},\bm{\Lambda}^{\star})$
    when $\alpha\in(\frac{1}{2},1)$; non-asymptotically speaking, the variance of this term is $\overline{\bm{\Lambda}}_{T}/T$ by definition, and $\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}=\bm{O}(T^{\alpha-1})$;
  3. 3.
    The term $I_{3}$ is related to both the samples $\{(\bm{A}_{t},\bm{b}_{t})\}_{1\leq t\leq T}$ and the intermediate iterations $\{\bm{\theta}_{t}\}_{0\leq t\leq T}$. Hence, we refer to this term as the iterative error. According to our analysis, this term converges by the rate of
    $\left\|\frac{1}{T}\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}-\bm{A})\bm{\Delta}_{i-1}\right\|_{2}\lesssim O(T^{-\frac{\alpha}{2}}).$
Comparing the magnitudes of the components.
Figure 8 illustrates the convergence of these terms in our $10,000$ independent trials when $d=3$ and $\alpha=\frac{2}{3}$. In the plot, the blue, orange and green lines represent the sample averages of $\|I_{1}\|_{2}^{2},\|I_{2}\|_{2}^{2}$ and $\|I_{3}\|_{2}^{2}$ respectively; the red line represents the evolvement of $\|\overline{\bm{\Delta}}_{T}\|_{2}^{2}$, and the black dotted line represents the asymptotic variance $\mathsf{Tr}(\bm{\Lambda}^{\star})/T$. As is shown in the graph, the initialization error $I_{1}$ decreases the fastest; when $T<10^{3}$, the iterative error $I_{2}$ is the dominant component in the estimation error, as the orange and red lines almost overlap. However, the slope of the orange line is steeper than that of the green line representing the sampling error $I_{3}$, an illustration that the iterative error converges faster as a function of $T$. In the mean time, the green line converges quickly to the black dotted line, verifying our claim that the sampling error corresponds to the asymptotic distribution. Finally, as $T$ becomes sufficiently large ($10^{5}$ in this example), the green, red, and black dotted lines overlap with each other, an illustration that the sampling error dominates the estimation error, which converges to its asymptotic distribution.
Figure 8: Decomposition of the TD estimation error through $10,000$ independent trials, using MDP with $d=3$ and stepsizes $\eta_{t}=\eta_{0}t^{-\frac{2}{3}}$
Difference between $\overline{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$.
Theorem A.11 features the difference between the "non-asymptotic variance" $\overline{\bm{\Lambda}}_{T}$ and the asymptotic variance $\bm{\Lambda}^{\star}$ as
$\overline{\bm{\Lambda}}_{T}-\bm{\Lambda}^{\star}\asymp\bm{O}(T^{\alpha-1}),$
which results in our conclusion that the difference between the non-asymptotic and asymptotic distributions of $\overline{\bm{\Delta}}_{T}$ is lower bounded by $O(T^{\alpha-1})$. In order to illustrate this difference in variance through numerical simulations, we use $\check{\bm{\Lambda}}_{T}$ to denote the sample mean (averaged over our $10,000$ independent trials) of
$\frac{1}{T}\left(\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right)\left(\sum_{i=1}^{T}\bm{Q}_{i}(\bm{A}_{i}\bm{\theta}^{\star}-\bm{b}_{i})\right)^{\top}.$
By definition, $\check{\bm{\Lambda}}_{T}$ is an unbiased estimator of $\overline{\bm{\Lambda}}_{T}$. Figure 9 demonstrates the difference between $\check{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$, measured by the Frobenius norm of their difference, as $T$ increases. It is clear from the plot that the difference converges slower when $\alpha$ is larger, and especially that it does not converge to 0 when $\alpha=1$. Notice that the plot is truncated before $T=10^{4}$ as when $T$ becomes larger than the number of trials, the randomness within $\check{\bm{\Lambda}}_{T}$ dominates the difference, making the plot uninformative.
Figure 9: The difference between $\check{\bm{\Lambda}}_{T}$ and $\bm{\Lambda}^{\star}$, for $d=3$ and different choices of $\alpha$
Appendix E EXTENSION TO GENERAL LINEAR STOCHASTIC APPROXIMATION
In this section, we explore on the conditions under which our theoretical analysis and statistical inference procedure can be extended to general Linear Stochastic Approximation (LSA) for the linear system
$\bm{A\theta}=\bm{b},$
where a sample matrix $\bm{A}_{t}$ and a sample vector $\bm{b}_{t}$ are observed $i.i.d.$ at each time $t$, with expectations being $\bm{A}$ and $\bm{b}$ respectively. Similar to TD learning, we also run
$\bm{\theta}_{t}=\bm{\theta}_{t-1}-\eta_{t}(\bm{A}_{t}\bm{\theta}_{t-1}-\bm{b}_{t}),$
with the same choice of stepsizes $\eta_{t}=\eta_{0}t^{-\alpha}$, $\alpha\in(\frac{1}{2},1)$.
We first make the following assumptions on the matrix $\bm{A}$:
Assumption E.1 (Stability).
There exists a $d\times d$ symmetric matrix $\bm{U}\succeq\bm{0}$ and a constant $\beta_{0}>0$, such that $\bm{AU}+\bm{UA}^{\top}\succeq\beta_{0}\bm{I}$;
Notice that in our setting of TD learning, this assumption is satisfied with $\bm{U}=\bm{I}$, and $\beta_{0}=2(1-\gamma)\lambda_{0}$; for general LSA, it is guaranteed as long as $\bm{A}$ is Hurwitz, i.e., all the eigenvalues of $\bm{A}$ has strictly negative real-parts. Under that condition, $\beta_{0}$ can be set to $1$, and $\bm{U}$ is the unique solution to the Lyapunov equation $\bm{AU}+\bm{UA}^{\top}=\bm{I}$. Meanwhile, we need the following constraints on the samples $\{(\bm{A}_{t},\bm{b}_{t})\}$:
Assumption E.2 (Sampling constraints).
With $\bm{U}$ specified as in Assumption E.1,
  1. 1.
    There exists a constant $c_{U}>0$, such that $\mathbb{E}[\bm{A}_{t}\bm{U}\bm{A}_{t}^{\top}]\preceq c_{U}(\bm{AU}+\bm{UA}^{\top})$;
  2. 2.
    There exists constants $c_{A}>0,c_{b}>0$, such that $\|\bm{A}_{t}\|_{\bm{U}}\leq c_{A}$ and $\|\bm{b}_{t}\|_{\bm{U}}\leq c_{b}$ almost surely.
Under our setting of TD learning, Assumption E.2 is satisfied with $c_{U}=1,c_{A}=2$ and $c_{b}=1$ by Lemma A.4. For general LSA with $\bm{A}$ being Hurwitz, as long as (2) is satisfied, (1) can also be guaranteed by setting $c_{U}=c_{A}\|\bm{U}\|$. As long as both Assumption E.2 and Assumption E.2 are satisfied, it is immediate that $\|\bm{A}_{t}\bm{\theta}^{\star}-\bm{b}_{t}\|_{\bm{U}}\leq c_{A}\|\bm{\theta}^{\star}\|_{\bm{U}}+c_{b}$ almost surely, and that for every $\bm{x}\in\mathbb{R}^{d}$,
$\|(\bm{I}-\eta\bm{A})\bm{x}\|_{\bm{U}}^{2}$ $\leq\mathbb{E}\|(\bm{I}-\eta\bm{A}_{t})\bm{x}\|_{\bm{U}}^{2}$
$=\|\bm{x}\|_{\bm{U}}^{2}-\eta\bm{x}^{\top}(\bm{AU}+\bm{UA}^{\top})\bm{x}+\eta^{2}\bm{x}^{\top}\mathbb{E}[\bm{A}_{t}\bm{U}\bm{A}_{t}^{\top}]\bm{x}$
$\leq\|\bm{x}\|_{\bm{U}}^{2}-(\eta-c_{U}\eta^{2})\bm{x}^{\top}(\bm{AU}+\bm{UA}^{\top})\bm{x};$
as long as
$\eta\leq\frac{1}{2c_{U}}=:\eta_{\max},$
this bound can be further elaborated as
$\|(\bm{I}-\eta\bm{A})\bm{x}\|_{\bm{U}}^{2}$ $\leq\mathbb{E}\|(\bm{I}-\eta\bm{A}_{t})\bm{x}\|_{\bm{U}}^{2}$
$\leq\|\bm{x}\|_{\bm{U}}^{2}-\frac{\eta}{2}\bm{x}^{\top}(\bm{AU}+\bm{UA}^{\top})\bm{x}$
$\leq\|\bm{x}\|_{\bm{U}}^{2}-\frac{\beta_{0}\eta}{2}\|\bm{x}\|_{2}^{2}\leq\left(1-\frac{\beta_{0}\eta}{2\|\bm{U}\|}\right)\|\bm{x}\|_{\bm{U}}^{2};$
Consequently, as long as $\eta\leq\eta_{\max}$,
$\|\bm{I}-\eta\bm{A}\|_{\bm{U}}\leq\sqrt{1-\frac{\beta_{0}\eta}{2\|\bm{U}\|}}\leq 1-\frac{\beta_{0}\eta}{4\|\bm{U}\|}.$
This is a direct generalization of Lemma A.4.
While the detailed derivations are well beyond the scope of this paper, our theoretical analysis in Theorems 3.1, 3.2, 3.3 can all be extended to general LSA with this generalized version of Lemma A.4. Specifically:
  1. 1.
    For every $\delta\in(0,1)$, there exists a proper choice of $\eta_{0}$, such that with probability at least $1-\delta$,
    $\|\overline{\bm{\Delta}}_{T}\|_{\bm{U}}\lesssim\sqrt{\frac{\mathsf{Tr}(\bm{U\Lambda}^{\star})}{T}\log\frac{1}{\delta}};$
  2. 2.
    As long as $\eta_{0}\leq\eta_{\max}$, it can be guaranteed that
    $d_{\mathsf{C}}(\sqrt{T}\overline{\bm{\Delta}}_{T},\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim C_{1}T^{-\frac{\alpha}{2}}+C_{2}T^{\alpha-1},$
    with coefficients $C_{1}$ and $C_{2}$ related to $\bm{U},\beta_{0},c_{U},c_{A},c_{b}$. Therefore, choosing $\alpha=\frac{2}{3}$ would still yield an $O(T^{-1/3})$ convergence rate;
  3. 3.
    With variance estimator $\widehat{\bm{\Lambda}}_{T}$ defined in the same way as (18),
    $d_{\mathsf{TV}}(\mathcal{N}(\bm{0},\widehat{\bm{\Lambda}}_{T}),\mathcal{N}(\bm{0},\bm{\Lambda}^{\star}))\lesssim O(T^{-\frac{1}{3}})$
    with probability at least $1-O(T^{-\frac{1}{3}})$;
  4. 4.
    Under the same statistical inference procedure as described in Section 3.4, the confidence region $\mathcal{C}_{\delta}$ still satisfies
    $\mathbb{P}(\bm{\theta}^{\star}\in\mathcal{C}_{\delta})$ $\geq 1-\delta-O(T^{-\frac{1}{3}}).$